Trucks arrive at an unloading terminal at an average rate of 120 per hour.

Trucks arrive at an unloading terminal at an average rate of 120 per hour. If the arrival of more than seven trucks in one minute will cause the terminal to be working at full capacity, what is the probability that the terminal will be working at full capacity?
Answer #1

expected number of arrivals in a minute =120/60 =2 arrivals/minute

probability that the terminal will be working at full capacity =P(more than 7 arrivals)=P(X>7)

=1-P(X<=7)

= 1-∑x=07 e-2*2x/x!= 1-0.5987 =0.4013

(

Note:

if using ti-84 use command :1-poissoncdf(7,7)
if using excel use commannd :1-poisson(7,7,true)